If A and B be two matrices then determinant(A*B)=determinant(A)*determinant(B).Is the statement true.If yes please prove it

Grade:Upto college level

1 Answers

pranav bhartia
14 years ago
it would be obviously true... bcause 12=3*4 and hence det(a*b)=det(a)*det(b)...